# MiTTenS: A Dataset for Evaluating Gender Mistranslation

### v2

@@ Abstract
Kevin Robinson   Sneha Kudugunta♢‡   Romina Stella   Sunipa Dev   Jasmijn Bastings Google Research   Google DeepMind   University of Washington {kevinrobinson,snehakudugunta,romistella,sunipadev,bastings}@google.com
@@ 1 Introduction
Recent advances have enabled general-purpose foundation models with powerful multilingual capabilities including translation Ouyang et al. (2022); OpenAI et al. (2023); Chung et al. (2022); Gemini Team Google (2023). These models can be used as building blocks in a wide range of products and applications, foregrounding other work on gender bias in natural language processing more broadly (Sun et al., 2019; Costa-jussà, 2019; Stanczak and Augenstein, 2021, i.a.).
@@ 4 Conclusion
This work aims to contribute to society and to human well-being by creating a new dataset and demonstrating how it can be used to avoid potential harms in translation systems. Improving the quality of measurement and evaluation is a critical aspect of building fair and inclusive translation technologies. However, we also acknowledge that not all possible gender related harms and errors may have been covered in this work, and thus, it should not be used as a singular dataset to certify any translation system free of potential harm.

### v3

@@ Abstract
Kevin Robinson   Sneha Kudugunta♢‡   Romina Stella   Sunipa Dev   Jasmijn Bastings Google DeepMind   Google Research   University of Washington {kevinrobinson,snehakudugunta,romistella,sunipadev,bastings}@google.com
@@ 1 Introduction
Recent advances have enabled general-purpose foundation models with powerful multilingual capabilities including translation Ouyang et al. (2022); OpenAI et al. (2023); Chung et al. (2022); Gemini Team Google (2023). These models can be used as building blocks in a wide range of products and applications, highlighting the importance of other work on gender bias in natural language processing more broadly (Sun et al., 2019; Costa-jussà, 2019; Stanczak and Augenstein, 2021, i.a.).
@@ 4 Conclusion
This work aims to contribute to society and to human well-being by creating a new dataset and demonstrating how it can be used to measure some potential harms in translation systems. Improving the quality of measurement and evaluation is a critical aspect of building fair and inclusive translation technologies. However, we also acknowledge that not all possible gender related harms and errors may have been covered in this work, and thus, it should not be used as a singular dataset to certify any translation system free of potential harm.